Find the absolute extreme values of the function on the interval.

-x^2+7x-10, 2 less than or equal to x less than or equal to 5
Answer #1

A ܐܣܝܐ f(0) =-22+72-10 125 X 5 5 f(W== (-2+72-10) = -23 +7 Fost the critical poind of f(x), f(x)=0 da > - 206+7=0 7 2727 >>